OpenStudy (akgorgeous):

Part A: Solve −np − 80 < 60 for n. Show your work. Part B: Solve 2a − 5d = 30 for d. Show your work.

OpenStudy (studygurl14):

A: just solve it like you would any equation

OpenStudy (studygurl14):

start by moving the -80 to the other side

OpenStudy (cp123):

add 80 to both sides of the equation. so you would get -np< 60 + 80. simplify. -np<140 you need to get -n alone in order to solve for it, so since it is being multiplied by p, you do the opposite by dividing both sides by p. -np/p< 140/p simplify: -n<140/p. thats it.

OpenStudy (cp123):

do the same thing for the next equation. subtract 2a from both sides and simplify: -5d= 30-2a. then divide each sides by -5 to isolate d. -5d/-5=30-2a/-5 simplify: d=30-2a/-5\[d=\frac{ 30-2a }{ -5 }\]